# Shen2018S18

### v1

@@ Abstract
We report discovery and spectroscopic follow-up of 21 astrometric binaries containing solar-type stars and dark companions with masses near $1.4\,M_{\odot}$. The simplest interpretation is that the companions are dormant neutron stars (NSs), though ultramassive white dwarfs (WDs) and tight WD+WD binaries cannot be fully excluded. We selected targets from Gaia DR3 astrometric binary solutions in which the luminous star is on the main sequence and the dynamically-implied mass of the unseen companion is (a) more than $1.25\,M_{\odot}$ and (b) too high to be any non-degenerate star or close binary. We obtained multi-epoch radial velocities (RVs) over a period of 650 days, spanning a majority of the orbits’ dynamic range in RV. The RVs broadly validate the astrometric solutions and significantly tighten constraints on companion masses. Several systems have companion masses that are unambiguously above the Chandrasekhar limit, while the rest have masses between 1.25 and 1.4 $M_{\odot}$. The orbits are significantly more eccentric at fixed period than those of typical WD + MS binaries, perhaps due to natal kicks. Metal-poor stars are overrepresented in the sample: three out of 21 objects (14%) have $\rm[Fe/H]\sim-1.5$ and are on halo orbits, compared to $\sim 0.5\%$ of the parent Gaia binary sample. The metal-poor stars are all strongly enhanced in lithium. The formation history of these objects is puzzling: it is unclear both how the binaries escaped a merger or dramatic orbital shrinkage when the NS progenitors were red supergiants, and how they remained bound when the NSs formed. Gaia has now discovered 3 black holes (BHs) in astrometric binaries with masses above $9\,M_{\odot}$, and 21 NSs with masses near $1.4\,M_{\odot}$. The lack of intermediate-mass objects in this sample is striking and significant, supporting the existence of a BH/NS mass bimodality over four orders of magnitude in orbital period.
@@ 1 Introduction
Most stars with initial masses $\gtrsim 8\,M_{\odot}$ leave behind neutron stars (NSs) when they die. Several thousand NSs are known in the Milky Way, a large majority of which are radio pulsars. Most ($>99\%$; Lorimer 2008) young pulsars are isolated. Yet, a large majority of the massive stars from which NSs form are in binaries, triples, and higher-order multiples (e.g. Sana et al., 2012; Moe & Di Stefano, 2017). The apparent mismatch in multiplicity properties of NSs and their progenitors hints that most massive binaries are destroyed during or prior to the formation of a NS. This destruction can come as a result of binary interaction leading to a stellar merger, or due to the binary becoming unbound during a supernova (SN).
All known companions to young pulsars are massive OB stars (Kaspi et al., 1996; Bassa et al., 2011; Shannon et al., 2014; Lyne et al., 2015; Andersen et al., 2023; van der Wateren et al., 2023). The same is true for all detached NS + main sequence (MS) binaries detected in X-rays (e.g. Reig, 2011). This likely reflects the fact that binaries containing a massive star are more likely to survive mass transfer and a supernova when the companion is also a massive star.
@@ 2 Sample selection
Calculating the mass of a star’s unseen companion from its astrometric solution requires an estimate of the mass of the star. In constructing their candidate sample, Shahaf et al. (2023b) used the IsocLum mass estimates calculated by Gaia Collaboration et al. (2023b). These estimates, which are available through the Gaia archive in the gaiadr3.binary_masses catalog, were inferred by comparison of the extinction-corrected colors and absolute magnitudes to a grid of PARSEC isochrones, with a prior that the metallicity is close to solar. These masses can thus be overestimated if the metallicities are subsolar, or underestimated if they are supersolar. We improve the mass estimates in Section 4.
Shahaf et al. (2023b) noted that their compact object binary candidate sample appeared to fall within two populations in the mass-eccentricty plane: one with a mean mass close to $0.6\,M_{\odot}$ and eccentricities below 0.2, and another with an apparent mean mass of $\sim 1.3\,M_{\odot}$ and a broad eccentricity distribution. It would be natural, they noted, to identify these two populations with WDs and NSs. Our follow-up has shown that the division is probably not so simple: some systems in the high-mass, high-eccentricity sample unambiguously host white dwarfs, as revealed by strong UV excess. Other objects have high eccentricities but companion masses below $1\,M_{\odot}$, which are implausibly low for a NS. Despite these caveats, our follow-up has shown that most of the low-eccentricity, lower-mass compact object candidates identified by Shahaf et al. (2023b) are WDs, and at least some of the higher-eccentricity, higher-mass companions are NSs. We return to this discussion in Section 6.2.
We initiated spectroscopic follow-up for a majority of the NS candidates identified by Shahaf et al. (2023b) in June 2022. We also carried out RV follow-up observations of some astrometrically-selected NS candidates not included in the Shahaf et al. (2023b) sample because their orbital periods slightly exceed 1000 days. Several of our candidates were also listed in other samples of compact object candidates from Gaia DR3, including the sample curated by Andrews et al. (2022). The observations presented here were obtained before April 2024, but our program is ongoing.
@@ 2.2 Completeness of the sample
We defer a full description of our follow-up program – including follow-up of suspected WDs and all objects that turned out to have spurious astrometric solutions – to future work. We suspect that our sample contains most of the NS-hosting binaries with $M_{2}\gtrsim 1.25\,M_{\odot}$ and astrometric solutions published in DR3. A handful of likely good candidates are not included because our RV follow-up has not yet covered enough of the orbits to confirm the astrometric solution with high confidence, and another handful were excluded because they are too faint ($G>15$) to be amenable for RV follow-up with the instruments at our disposal. Our sample does not contain any low-mass NSs with $M<1.25\,M_{\odot}$, a mass limit below which some NSs likely do exist (Ferdman et al., 2014; Martinez et al., 2015). As we discuss throughout the paper, it becomes increasingly challenging to distinguish between NSs and massive WDs at lower masses.
@@ 3.1 FEROS
We obtained 128 spectra with the Fiberfed Extended Range Optical Spectrograph (FEROS; Kaufer et al., 1999) on the 2.2m ESO/MPG telescope at La Silla Observatory (programs P109.A-9001, P110.A-9014, P111.A-9003, P112.A-6010, and P113.26XB). Some observations used $2\times 2$ binning to reduce readout noise at the expense of spectral resolution; the remainder used $1\times 1$ binning. The resulting spectra have resolution $R\approx 40,000$ ($2\times 2$ binning) and $R\approx 50,000$ ($1\times 1$ binning). Exposure times ranged from 1200 to 3600 seconds, depending on source brightness. We reduced the data using the CERES pipeline (Brahm et al., 2017), which performs bias-subtraction, flat fielding, wavelength calibration, and optimal extraction. The pipeline measures and corrects for small shifts in the wavelength solution during the course a night via simultaneous observations of a ThAr lamp with a second fiber.
@@ 3.2 TRES
We obtained 146 spectra using the Tillinghast Reflector Echelle Spectrograph (TRES; Fűrész, 2008) mounted on the 1.5 m Tillinghast Reflector telescope at the Fred Lawrence Whipple Observatory (FLWO) on Mount Hopkins, Arizona. TRES is a fibrefed echelle spectrograph with a wavelength range of 390–910 nm and a resolving power of $R\sim 44,000$. The spectra were extracted as described in Buchhave et al. (2010).
@@ 3.3 MIKE
We obtained 7 spectra with the Magellan Inamori Kyocera Echelle (MIKE) spectrograph on the Magellan Clay telescope at Las Campanas Observatory (Bernstein et al., 2003). We used the 0.5” slit and exposure times ranging from 600 to 1200 seconds, yielding spectral resolution $R\sim 40,000$ on the blue side and $R\sim 55,000$ on the red side. The typical SNR at 5800 Å  was $\sim 30$ and the total wavelength coverage was $\sim 3330-9680$. We reduced the spectra with the MIKE Pipeline within CarPy (Kelson et al., 2000; Kelson, 2003) and subsequently flux-calibrated them using observations of a standard star. We merged the orders into a single spectrum, weighting by inverse variance in the overlap regions. We use the MIKE spectra (as well as the HIRES and PEPSI spectra described below) primarily for measuring stellar parameters and metallicities rather than RVs.
@@ 3.8 RVs and offsets
As described by El-Badry et al. (2024), we fit for a single global RV offset between TRES and FEROS. We found an offset of $0.16\,\rm km\,s^{-1}$, in the sense that ${\rm RV}_{{\rm FEROS}}={\rm RV}_{{\rm TRES}}-0.16\,{\rm km\,s^{-1}}$. This offset has already been applied to our reported RVs. Because we generally only obtained one spectrum per target for MIKE, HIRES, PEPSI, and ESI, we did not fit for offsets for these instruments, but instead adopted a generous 0.5 $\rm km\,s^{-1}$ minimum uncertainty for all the RVs measured with these instruments. All the RVs are listed in Appendix D and provided in machine-readable form as supplementary material.
@@ 3.9.3 Gaia XP metallicities
Metallicity SED fit Name [Fe/H], High-res Instrument [M/H], Gaia XP $E(B-V)$ $R\,[R_{\odot}]$ $T_{\rm eff}$ [kK] $M\,[M_{\odot}]$ $[\rm Fe/H]_{\rm init}$ J0553-1349 $0.12\pm 0.07$ MIKE $0.04\pm 0.03$ $0.06\pm 0.02$ $0.988\pm 0.015$ $5.59\pm 0.07$ $0.96\pm 0.04$ $0.14\pm 0.04$ J2057-4742 $0.13\pm 0.08$ MIKE $0.22\pm 0.06$ $0.01\pm 0.01$ $0.941\pm 0.010$ $5.834\pm 0.027$ $1.05\pm 0.03$ $0.096\pm 0.030$ J1553-6846 $0.13\pm 0.10$ MIKE $0.16\pm 0.06$ $0.05\pm 0.02$ $0.971\pm 0.013$ $5.88\pm 0.05$ $1.082\pm 0.03$ $0.141\pm 0.034$ J2102+3703 $-0.37\pm 0.08$ HIRES $-0.30\pm 0.05$ $0.06\pm 0.01$ $0.947\pm 0.010$ $6.27\pm 0.06$ $1.035\pm 0.03$ $-0.31\pm 0.04$ J0742-4749 $-0.13\pm 0.12$ MIKE $-0.11\pm 0.07$ $0.16\pm 0.02$ $1.259\pm 0.027$ $5.68\pm 0.07$ $0.958\pm 0.03$ $0.10\pm 0.05$ J0152-2049 $-1.28\pm 0.08$ TRES $-1.31\pm 0.17$ $0.04\pm 0.00$ $1.169\pm 0.013$ $6.47\pm 0.04$ $0.850\pm 0.03$ $-1.243\pm 0.015$ J0003-5604 $-0.09\pm 0.08$ FEROS $-0.03\pm 0.05$ $00.00\pm 0$ 0.00 $0.752\pm 0.006$ $4.885\pm 0.013$ $0.794\pm 0.03$ $-0.121\pm 0.024$ J1733+5808 $0.17\pm 0.10$ TRES $0.15\pm 0.09$ $0.05\pm 0.02$ $1.081\pm 0.019$ $6.01\pm 0.08$ $1.17\pm 0.05$ $0.23\pm 0.04$ J1150-2203 $0.42\pm 0.07$ MIKE $0.39\pm 0.05$ $0.05\pm 0.01$ $1.495\pm 0.019$ $5.834\pm 0.019$ $1.193\pm 0.03$ $0.297\pm 0.004$ J1449+6919 $-0.65\pm 0.09$ HIRES $-0.47\pm 0.06$ $0.05\pm 0.02$ $1.060\pm 0.024$ $6.03\pm 0.14$ $0.91\pm 0.05$ $-0.33\pm 0.05$ J0217-7541 $0.14\pm 0.10$ FEROS $0.15\pm 0.07$ $0.02\pm 0.01$ $1.238\pm 0.018$ $5.58\pm 0.05$ $0.988\pm 0.03$ $0.272\pm 0.025$ J0639-3655 $0.04\pm 0.10$ MIKE $0.09\pm 0.04$ $0.09\pm 0.03$ $1.43\pm 0.04$ $6.48\pm 0.21$ $1.32\pm 0.06$ $0.17\pm 0.05$ J1739+4502 $-1.82\pm 0.10$ MIKE $-1.46\pm 0.23$ $0.06\pm 0.00$ $1.424\pm 0.016$ $6.289\pm 0.032$ $0.78\pm 0.03$ $-1.58\pm 0.04$ J0036-0932 $-0.42\pm 0.10$ HIRES $-0.25\pm 0.05$ $0.03\pm 0.02$ $1.32\pm 0.04$ $5.84\pm 0.09$ $0.94\pm 0.04$ $-0.10\pm 0.05$ J1432-1021 $-1.29\pm 0.11$ MIKE $-1.48\pm 0.26$ $0.10\pm 0.00$ $1.491\pm 0.029$ $6.049\pm 0.023$ $0.79\pm 0.03$ $-1.239\pm 0.014$ J1048+6547 $-0.14\pm 0.11$ HIRES $-0.19\pm 0.08$ $0.02\pm 0.02$ $1.15\pm 0.04$ $5.96\pm 0.10$ $1.00\pm 0.05$ $-0.03\pm 0.06$ J2145+2837 $-0.03\pm 0.06$ HIRES $0.21\pm 0.04$ $0.01\pm 0.01$ $0.846\pm 0.005$ $5.500\pm 0.023$ $0.948\pm 0.03$ $0.022\pm 0.023$ J2244-2236 $0.06\pm 0.08$ TRES $0.03\pm 0.07$ $0.00\pm 0.01$ $0.881\pm 0.009$ $5.892\pm 0.026$ $0.990\pm 0.03$ $-0.127\pm 0.034$ J0824+5254 $0.27\pm 0.10$ TRES $0.13\pm 0.07$ $0.04\pm 0.02$ $0.997\pm 0.013$ $5.89\pm 0.05$ $1.104\pm 0.03$ $0.19\pm 0.04$ J0230+5950 $0.15\pm 0.08$ HIRES $0.11\pm 0.06$ $0.21\pm 0.01$ $1.001\pm 0.008$ $5.87\pm 0.04$ $1.114\pm 0.03$ $0.230\pm 0.034$ J0634+6256 $-0.03\pm 0.11$ TRES $-0.23\pm 0.11$ $0.12\pm 0.02$ $1.68\pm 0.07$ $6.05\pm 0.12$ $1.17\pm 0.06$ $0.08\pm 0.06$

### v2

@@ Abstract
We report discovery and spectroscopic follow-up of 21 astrometric binaries containing solar-type stars and dark companions with masses near $1.4\,M_{\odot}$. The simplest interpretation is that the companions are dormant neutron stars (NSs), though ultramassive white dwarfs (WDs) and tight WD+WD binaries cannot be fully excluded. We selected targets from Gaia DR3 astrometric binary solutions in which the luminous star is on the main sequence and the dynamically-implied mass of the unseen companion is (a) more than $1.25\,M_{\odot}$ and (b) too high to be any non-degenerate star or close binary. We obtained multi-epoch radial velocities (RVs) over a period of 700 days, spanning a majority of the orbits’ dynamic range in RV. The RVs broadly validate the astrometric solutions and significantly tighten constraints on companion masses. Several systems have companion masses that are unambiguously above the Chandrasekhar limit, while the rest have masses between 1.25 and 1.4 $M_{\odot}$. The orbits are significantly more eccentric at fixed period than those of typical WD + MS binaries, perhaps due to natal kicks. Metal-poor stars are overrepresented in the sample: three out of 21 objects (14%) have $\rm[Fe/H]\sim-1.5$ and are on halo orbits, compared to $\sim 0.5\%$ of the parent Gaia binary sample. The metal-poor stars are all strongly enhanced in lithium. The formation history of these objects is puzzling: it is unclear both how the binaries escaped a merger or dramatic orbital shrinkage when the NS progenitors were red supergiants, and how they remained bound when the NSs formed. Gaia has now discovered 3 black holes (BHs) in astrometric binaries with masses above $9\,M_{\odot}$, and 21 NSs with masses near $1.4\,M_{\odot}$. The lack of intermediate-mass objects in this sample is striking and significant, supporting the existence of a BH/NS mass bimodality over four orders of magnitude in orbital period.
@@ 1 Introduction
Most stars with initial masses $\gtrsim 8\,M_{\odot}$ leave behind neutron stars (NSs) when they die. Several thousand NSs are known in the Milky Way, a large majority of which are radio pulsars. Most ($>99\%$; Lorimer 2008) young pulsars are isolated. Yet, a large majority of the massive stars from which NSs form are in binaries, triples, and higher-order multiples (e.g. Sana et al., 2012; Moe & Di Stefano, 2017). The apparent mismatch in multiplicity properties of NSs and their progenitors hints that most massive binaries are destroyed during or prior to the formation of a NS. This destruction can come as a result of binary interaction leading to a stellar merger, or due to the binary becoming unbound during a supernova (SN), which can impart a kick of order $250\,\rm km\,s^{-1}$ on the newborn NS (e.g. Hobbs et al., 2005; Faucher-Giguère & Kaspi, 2006).
All known companions to young pulsars are massive OB stars (Kaspi et al., 1996; Bassa et al., 2011; Shannon et al., 2014; Lyne et al., 2015; Andersen et al., 2023; van der Wateren et al., 2023). The same is true for all detached NS + main sequence (MS) binaries detected in X-rays (e.g. Reig, 2011). This likely reflects the fact that binaries containing a massive star are more likely to survive mass transfer and a SN when the companion is also a massive star.
@@ 2 Sample selection
Calculating the mass of a star’s unseen companion from its astrometric solution requires an estimate of the mass of the star. In constructing their candidate sample, Shahaf et al. (2023b) used the IsocLum mass estimates calculated by Gaia Collaboration et al. (2023b). These estimates, which are available through the Gaia archive in the gaiadr3.binary_masses catalog, were inferred by comparison of the extinction-corrected colors and absolute magnitudes to a grid of PARSEC isochrones, with a prior that the metallicity is close to solar. These masses can thus be overestimated if the metallicities are subsolar, or underestimated if they are supersolar. Their validity is also contingent on the validity of the extinction estimates and on the assumption that a single star contributes to the observed photometry. We improve the mass estimates in Section 4.
Shahaf et al. (2023b) noted that their compact object binary candidate sample appeared to fall within two populations in the mass-eccentricty plane: one with a mean mass close to $0.6\,M_{\odot}$ and eccentricities below 0.2, and another with an apparent mean mass of $\sim 1.3\,M_{\odot}$ and a broad eccentricity distribution. It would be natural, they noted, to identify these two populations with WDs and NSs. Our follow-up has shown that the division is probably not so simple: some systems in the high-mass, high-eccentricity sample unambiguously host white dwarfs, as revealed by strong UV excess (e.g. Ganguly et al., 2023). Other objects have high eccentricities but companion masses below $1\,M_{\odot}$, which are implausibly low for a NS. Despite these caveats, our follow-up has shown that most of the low-eccentricity, lower-mass compact object candidates identified by Shahaf et al. (2023b) are WDs, and at least some of the higher-eccentricity, higher-mass companions are NSs. We return to this discussion in Section 6.2.
We initiated spectroscopic follow-up for a majority of the NS candidates identified by Shahaf et al. (2023b) in June 2022. We also carried out RV follow-up observations of some astrometrically-selected NS candidates not included in the Shahaf et al. (2023b) sample because their orbital periods slightly exceed 1000 days. Several of our candidates were also listed in other samples of compact object candidates from Gaia DR3, including the sample curated by Andrews et al. (2022). The observations presented here were obtained before June 2024, but our program is ongoing.
@@ 2.2 Completeness of the sample
Table 3 in Appendix A provides a summary of our RV follow-up and our current assessment of the viability of all candidates from the Shahaf et al. (2023b) and Andrews et al. (2022) samples. We defer a full description of our follow-up program – including RVs of suspected WDs and all objects that turned out to have spurious astrometric solutions – to future work. We suspect that our sample contains most of the NS-hosting binaries with $M_{2}\gtrsim 1.25\,M_{\odot}$ and astrometric solutions published in DR3. A handful of likely good candidates are not included because our RV follow-up has not yet covered enough of the orbits to confirm the astrometric solution with high confidence, and another handful were excluded because they are too faint ($G>15$) to be amenable for RV follow-up with the instruments at our disposal. Our sample does not contain any low-mass NSs with $M<1.25\,M_{\odot}$, a mass limit below which some NSs likely do exist (Ferdman et al., 2014; Martinez et al., 2015). As we discuss throughout the paper, it becomes increasingly challenging to distinguish between NSs and massive WDs at lower masses.
@@ 3.1 FEROS
We obtained 129 spectra with the Fiberfed Extended Range Optical Spectrograph (FEROS; Kaufer et al., 1999) on the 2.2m ESO/MPG telescope at La Silla Observatory (programs P109.A-9001, P110.A-9014, P111.A-9003, P112.A-6010, and P113.26XB). Some observations used $2\times 2$ binning to reduce readout noise at the expense of spectral resolution; the remainder used $1\times 1$ binning. The resulting spectra have resolution $R\approx 40,000$ ($2\times 2$ binning) and $R\approx 50,000$ ($1\times 1$ binning). Exposure times ranged from 1200 to 3600 seconds, depending on source brightness. We reduced the data using the CERES pipeline (Brahm et al., 2017), which performs bias-subtraction, flat fielding, wavelength calibration, and optimal extraction. The pipeline measures and corrects for small shifts in the wavelength solution during the course a night via simultaneous observations of a ThAr lamp with a second fiber.
@@ 3.2 TRES
We obtained 155 spectra using the Tillinghast Reflector Echelle Spectrograph (TRES; Fűrész, 2008) mounted on the 1.5 m Tillinghast Reflector telescope at the Fred Lawrence Whipple Observatory (FLWO) on Mount Hopkins, Arizona. TRES is a fibrefed echelle spectrograph with a wavelength range of 390–910 nm and a resolving power of $R\sim 44,000$. The spectra were extracted as described in Buchhave et al. (2010).
@@ 3.3 MIKE
We obtained 7 spectra with the Magellan Inamori Kyocera Echelle (MIKE) spectrograph on the Magellan Clay telescope at Las Campanas Observatory (Bernstein et al., 2003). We used the 0.5” slit and exposure times ranging from 600 to 1200 seconds, yielding spectral resolution $R\sim 40,000$ on the blue side and $R\sim 55,000$ on the red side. The typical SNR at 5800 Å  was $\sim 30$ and the total wavelength coverage was $\sim 3330-9680$. We reduced the spectra with the MIKE Pipeline within CarPy (Kelson et al., 2000; Kelson, 2003) and subsequently flux-calibrated them using observations of a standard star. We merged the orders into a single spectrum, weighting by inverse variance in the overlap regions.
@@ 3.8 RVs and offsets
As described by El-Badry et al. (2024), we fit for a single global RV offset between TRES and FEROS. We found an offset of $0.16\,\rm km\,s^{-1}$, in the sense that ${\rm RV}_{{\rm FEROS}}={\rm RV}_{{\rm TRES}}-0.16\,{\rm km\,s^{-1}}$. This offset has already been applied to our reported RVs. Because we generally only obtained one spectrum per target for MIKE, HIRES, PEPSI, and ESI, we did not fit for offsets for these instruments, but instead adopted a conservative 0.5 $\rm km\,s^{-1}$ minimum uncertainty for all the RVs measured with these instruments. All the RVs are listed in Appendix D and provided in machine-readable form as supplementary material.
@@ 3.9.3 Gaia XP metallicities
Metallicity SED fit Name [Fe/H], High-res Instrument [M/H], Gaia XP $E(B-V)$ $R\,[R_{\odot}]$ $T_{\rm eff}$ [kK] $M\,[M_{\odot}]$ $[\rm Fe/H]_{\rm init}$ J0553-1349 $0.12\pm 0.07$ MIKE $0.04\pm 0.03$ $0.06\pm 0.02$ $0.988\pm 0.015$ $5.59\pm 0.07$ $0.98\pm 0.06$ $0.14\pm 0.04$ J2057-4742 $0.13\pm 0.08$ MIKE $0.22\pm 0.06$ $0.01\pm 0.01$ $0.941\pm 0.010$ $5.834\pm 0.027$ $1.05\pm 0.03$ $0.096\pm 0.030$ J1553-6846 $0.13\pm 0.10$ MIKE $0.16\pm 0.06$ $0.05\pm 0.02$ $0.971\pm 0.013$ $5.88\pm 0.05$ $1.04\pm 0.05$ $0.141\pm 0.034$ J2102+3703 $-0.37\pm 0.08$ HIRES $-0.30\pm 0.05$ $0.06\pm 0.01$ $0.947\pm 0.010$ $6.27\pm 0.06$ $1.035\pm 0.03$ $-0.31\pm 0.04$ J0742-4749 $-0.13\pm 0.12$ MIKE $-0.11\pm 0.07$ $0.16\pm 0.02$ $1.259\pm 0.027$ $5.68\pm 0.07$ $0.90\pm 0.05$ $0.10\pm 0.05$ J0152-2049 $-1.28\pm 0.08$ TRES $-1.31\pm 0.17$ $0.04\pm 0.00$ $1.169\pm 0.013$ $6.47\pm 0.04$ $0.780\pm 0.03$ $-1.243\pm 0.015$ J0003-5604 $-0.09\pm 0.08$ FEROS $-0.03\pm 0.05$ $00.00\pm 0$ 0.00 $0.752\pm 0.006$ $4.885\pm 0.013$ $0.794\pm 0.03$ $-0.121\pm 0.024$ J1733+5808 $0.17\pm 0.10$ TRES $0.15\pm 0.09$ $0.05\pm 0.02$ $1.081\pm 0.019$ $6.01\pm 0.08$ $1.17\pm 0.05$ $0.23\pm 0.04$ J1150-2203 $0.42\pm 0.07$ MIKE $0.39\pm 0.05$ $0.05\pm 0.01$ $1.495\pm 0.019$ $5.834\pm 0.019$ $1.18\pm 0.06$ $0.297\pm 0.004$ J1449+6919 $-0.65\pm 0.09$ HIRES $-0.47\pm 0.06$ $0.05\pm 0.02$ $1.060\pm 0.024$ $6.03\pm 0.14$ $0.91\pm 0.05$ $-0.33\pm 0.05$ J0217-7541 $0.14\pm 0.10$ FEROS $0.15\pm 0.07$ $0.02\pm 0.01$ $1.238\pm 0.018$ $5.58\pm 0.05$ $0.99\pm 0.03$ $0.272\pm 0.025$ J0639-3655 $0.04\pm 0.10$ MIKE $0.09\pm 0.04$ $0.09\pm 0.03$ $1.43\pm 0.04$ $6.48\pm 0.21$ $1.32\pm 0.06$ $0.17\pm 0.05$ J1739+4502 $-1.82\pm 0.10$ MIKE $-1.46\pm 0.23$ $0.06\pm 0.00$ $1.424\pm 0.016$ $6.289\pm 0.032$ $0.78\pm 0.03$ $-1.58\pm 0.04$ J0036-0932 $-0.42\pm 0.10$ HIRES $-0.25\pm 0.05$ $0.03\pm 0.02$ $1.32\pm 0.04$ $5.84\pm 0.09$ $0.94\pm 0.04$ $-0.10\pm 0.05$ J1432-1021 $-1.29\pm 0.11$ MIKE $-1.48\pm 0.26$ $0.10\pm 0.00$ $1.491\pm 0.029$ $6.049\pm 0.023$ $0.79\pm 0.03$ $-1.239\pm 0.014$ J1048+6547 $-0.14\pm 0.11$ HIRES $-0.19\pm 0.08$ $0.02\pm 0.02$ $1.15\pm 0.04$ $5.96\pm 0.10$ $1.00\pm 0.05$ $-0.03\pm 0.06$ J2145+2837 $-0.03\pm 0.06$ HIRES $0.21\pm 0.04$ $0.01\pm 0.01$ $0.846\pm 0.005$ $5.500\pm 0.023$ $0.948\pm 0.03$ $0.022\pm 0.023$ J2244-2236 $0.06\pm 0.08$ TRES $0.03\pm 0.07$ $0.00\pm 0.01$ $0.881\pm 0.009$ $5.892\pm 0.026$ $0.990\pm 0.03$ $-0.127\pm 0.034$ J0824+5254 $0.27\pm 0.10$ TRES $0.13\pm 0.07$ $0.04\pm 0.02$ $0.997\pm 0.013$ $5.89\pm 0.05$ $1.104\pm 0.03$ $0.19\pm 0.04$ J0230+5950 $0.15\pm 0.08$ HIRES $0.11\pm 0.06$ $0.21\pm 0.01$ $1.001\pm 0.008$ $5.87\pm 0.04$ $1.114\pm 0.03$ $0.230\pm 0.034$ J0634+6256 $-0.03\pm 0.11$ TRES $-0.23\pm 0.11$ $0.12\pm 0.02$ $1.68\pm 0.07$ $6.05\pm 0.12$ $1.17\pm 0.06$ $0.08\pm 0.06$